Use Euler's method with step size 0.2 to estimate y(1), where y(x) is the solution of the initial-value problem
y'=2x+y^2,   y(0)=−1.

y(1)= .
Answer #1

we have, y = 2x+y?, with y(0) -- ilet f(x, y) = 2x+y2 & 20 = 0 & Yo=-1 It is given that, step size ,i-e., h=0.2 ..Xo=0 2.2,=Now, Ya y (22) y (0-4)= 4, thif(24, Y) - 60.8) +0.2) f (0.2,-0.8) =(-0.8) + (0.2) (2x0.2+(-0.834 = 60.8) +(0:2) (1.06) Yz= y(Now, Y = 4(25)=y (1) = 44th.f(244, Ye) = 60.09572) +(0.2)F(0.8,-0.0957) = (-0.09572)+(0.2) (2x0.8+ (-0.09572) = 60.09572)+(0.